OpenStudy (anonymous):

You bought a new car, but don’t remember the original purchase price. The monthly payments are $379.50 for 36 months. The salesperson said the simple interest rate for the period of the loan was 6%. You were also allowed $2500 for your old car. Find the original purchase price of the new car. Plz show me your work so I can know how you got the answer.

OpenStudy (hoblos):

379.5 * 36 = 13662 13662 - 6%= 12842.28 12842.28 + 2500=$15342.28

OpenStudy (anonymous):

ty, I see what I was doing wrong now.

OpenStudy (hoblos):

:D